# Optimal Clustering with Dependent Costs in Bayesian Networks

### v4

@@ Abstract
journal: Expert Systems with Applications
[…]
organization=Centre for Data Science, School of Mathemmatical Sciences, Queensland University of Technology,addressline=2 George Street, city=Brisbane, postcode=4000, state=Queensland, country=Australia, email=. p.wu@qut.edu.au
[…]
organization=Institute of Applied Mathematics and Information Technology, Italian National Research Councily,addressline=Via Alfonso Corti 12, city=Milano, postcode=I-20133, state=, country=Italy, email=. fabrizio@mi.imati.cnr.it
[…]
organization=Centre for Data Science, School of Mathemmatical Sciences, Queensland University of Technology,addressline=2 George Street, city=Brisbane, postcode=4000, state=Queensland, country=Australia, email=. k.mengersen@qut.edu.au
{highlights}
Clustering of nodes in Bayesian Networks can improve computation and model learning
Cost dependency makes global optimisation challenging
Novel algorithm Dependent Cluster MAPping (DCMAP) finds least cost cluster mappings
An equally informed algorithm cannot use fewer iterations than DCMAP
Demonstrated on seagrass DBN to improve computation, potential for minimising entropy
@@ 1 Introduction
Bayesian Networks (BNs) and related methods including Dynamics BNs (DBNs) are a form of graphical model that are represented as Directed Acyclic Graphs (DAGs), where relationships between nodes (factors) are shown as arcs and quantified with conditional probabilities (Pearl, 1988). They have been applied widely across domains including medicine, environment, economics and social sciences to provide explanatory decision support, integrating data and even expert knowledge (Wu et al., 2017; Kitson et al., 2023). However, one of the key challenges associated with inference and learning of BNs and DBNs is computation, which is NP-hard (Park and Darwiche, 2004; Murphy, 2002). To this end, different approaches to clustering of BN nodes have been used in the context of supporting more computationally efficient inference of marginal distributions given evidence, and model learning, to update the structure of the graph and/or the conditional probabilities associated with each node (Daly et al., 2011; Lauritzen and Spiegelhalter, 1988; Kojima et al., 2010; Lu et al., 2021).
Optimal clustering of a BN seeks to minimise some cost or objective function associated with the mapping of BN nodes to clusters. Typically, each node is uniquely mapped to a cluster and the total cost is the sum of costs associated with each cluster. However, a key challenge is cost dependency. Consider a BN $H(\boldsymbol{X},\boldsymbol{A})$ with nodes $X_{i}\in\boldsymbol{X}$ and arcs (or edges) $\boldsymbol{A}$ that encode conditional dependence between parent and child nodes. Fundamentally, BN inference is underpinned by Bayes theorem where the joint probability distribution over all nodes $\boldsymbol{X}$ and given observations (i.e. evidence $\boldsymbol{E}$) is (Lauritzen and Spiegelhalter, 1988):
@@
where $P(X_{j}|Par(X_{j}))$ is the conditional probability of node $X_{j}$ given its parents, and evidence about a node $X_{j}$ is notated as $\delta(X_{j})\in[0,1]$, $\sum\delta(X_{j})=1$. The joint probability distribution arises from the multiplication of conditional probability distributions, each of which is potentially multi-dimensional, comprising a node (or variable) $X_{j}$ and one or more parent nodes $Par(X_{j})$. Hence, intuitively, if nodes were clustered then the outcomes of inference at a given cluster are dependent on clusters connected by at least one arc via multiplication as per equation (1). This propagation of partial inference outcomes between clusters has been employed in approaches to BN and DBN inference such as Lauritzen and Spiegelhalter (1988) and Wu et al. (2018). The dependency of the local cluster cost on the mapping of connected clusters is referred to as cost dependency, noting that the objective is to minimise the global cost which is a combination (typically additive) of local costs. However, finding optimal clusters (i.e. partitions) is NP-hard in general (Buluç et al., 2016) and cost dependency exacerbates this complexity.
The clustering problem can be formalised as one of assigning a unique integer cluster label $u(X_{j})=k,k=1,\ldots,m$ to each node $X_{j}\in\boldsymbol{X}$ of BN $H(\boldsymbol{X},\boldsymbol{A})$; the total number of clusters $m$ is determined through optimisation. Let $G\bigl{(}H(\boldsymbol{X},\boldsymbol{A}),u(\boldsymbol{X})\bigr{)}$ be the total cost to be minimised through cluster mapping and $G^{\ast}\bigl{(}H(\boldsymbol{X},\boldsymbol{A})\bigr{)}$ the optimal cost, and assume that the total cost is the sum of individual cluster costs $G_{k}$:
@@ 1.1 Case Study
We use as a case study a DBN of a seagrass ecosystem (Wu et al., 2018). Seagrasses are a critical primary habitat for fish and many endangered species including the dugong and green turtle, and contribute USD\$1.9 trillion per annum through nutrient uptake and cycling. However, better management of human stressors such as dredging requires the modelling of cumulative effects arising from interactions between biological, ecological and environmental factors. The non-homogeneous DBN itself is complex and computationally intensive, and inference is infeasible within a standard (MCMC) framework for real-world decision support contexts. Furthermore, understanding areas of uncertainty in the system and designing data collection strategies to mitigate uncertainty is of utmost interest in managing such complex systems. This uncertainty can be characterised by entropy with clusters corresponding to scenarios of evidence (i.e. observed data) (Section 5.2). We focus on computation as a cost function $G$ to demonstrate how the proposed algorithm can find clusters to minimise the computational complexity of inference, and discuss how our algorithm could also be used to minimise other cost functions such as entropy.
@@ 1.2 Existing Work
Clustering of BNs has been employed in a wide range of contexts, evolving from a focus on computation of marginal distributions given evidence such as the seminal work of Lauritzen and Spiegelhalter (1988) to predominantly causal discovery and graph learning (Kitson et al., 2023; Lu et al., 2021) in recent times. Although the approach of Lauritzen and Spiegelhalter (1988) and related approaches such as Kschischang et al. (2001) (see Daly et al. (2011) for a review) employ clustering, they are based on a single mapping of nodes to clusters (referred to as cliques) using heuristics. They do not evaluate different cluster mappings on an outcome score. These challenges are addressed in more recent literature through three main approaches: (i) constraint based, that evaluate conditional independence structures i.e. if an arc exists or not, then its direction, (ii) scoring, often with heuristics to evaluate goodness of fit of a proposed network, and (iii) hybrid, combining constraints with heuristics (Huang and Zhou, 2022) .
@@
A number of approaches use clusters as a structure to restrict or constrain the local search space in a graph learning framework such as the seminal work of Friedman et al. (1999). They define clusters in a similar vein to clique-trees (Lauritzen and Spiegelhalter, 1988) in a “restrict and maximise” framework to optimise a mutual-information based score of the learned network. Kojima et al. (2010) advance this approach, effectively relaxing the cluster tree requirement in Friedman et al. (1999) to a cluster graph to help address the issue of large clusters, which make local optimisation computationally infeasible. They used ancestral constraints to help limit the search space of clusters and the algorithm was shown to outperform greedy algorithms including Max-Min Hill Climbing (MMHC) and Hill Climbing (HC) with Tabu search. However, computational feasibility becomes a problem again when there are too many edges between clusters (Kojima et al., 2010).
@@
Lu et al. (2021) also adapt the work of Friedman et al. (1999) for undirected graphs, using clusters based on proposed parents and children of a node $X$ to constrain the search in an iterative framework based on A*, a heuristic-informed search algorithm. They show empirically that for small networks, exhaustive and in some cases optimal methods such as A* and SAT (Boolean SATisfiability) outperform greedy methods such as PC (Peter and Clark), which can suffer from convergence to local optimal, and their adaptation of A* performs as well as greedy methods in larger networks. However, their approach is exponential in complexity with neighbourhood size (i.e. number of parents).
@@
Zhang et al. (2018) present an alternate approach in which clusters are determined through agglomerative clustering using Pearson correlation as the distance metric; however, little is provided in terms of algorithm performance and its convergence characteristics. Huang and Zhou (2022) and Gu and Zhou (2020) both develop comprehensive approaches that use similar agglomerative clustering based on mutual information and correlation as the distance metric, respectively. The former learns skeletons (an undirected version of the graph) for each cluster using their variant of the PC algorithm and log-likelihood ratio tests of conditional independence. Gu and Zhou (2020) use a score-based algorithm to learn partial DAGs for each cluster, which are then fused in the final step proposing edges between clusters. Huang and Zhou (2022) combine clusters with automated parameter tuning for turning skeletons into partial then fully directed acyclic graphs and hybrid greedy initialisation, obtaining empirical improvements in computational efficiency and more accurate DAGs. However, in both cases, the performance is dependent on the quality and size of the clusters.
@@
Clearly, clustering has demonstrated utility to improve computational feasibility in learning BNs of varying degrees of optimality given a scoring function (i.e. cost function). However, all of the reviewed methods above focused on scores that changed given changes to the BN vis a vis inclusion or exclusion of edges, rather than scores that changed given different mappings of nodes to clusters. This paper focuses on the latter task of explicit optimisation with respect to clusters, to develop a more generalisable algorithm that works with arbitrary non-negative objective functions. Nevertheless, the literature does highlight the utility of dynamic programming and clustering for optimisation (Friedman et al., 1999; Kojima et al., 2010; Lu et al., 2021).
@@
Note that there also exists substantial literature in graph partitioning such as repeated graph bisection, Kernighan and Lin algorithm, spectral bisection and multilevel partitioning (Buluç et al., 2016; Bader et al., 2013). Zheng et al. (2020), for instance, demonstrate how dynamic programming can be used to optimally and efficiently cluster nodes in the absence of cost dependency, applied to convolutional neural networks. However, the problem context for these approaches typically involves partitioning a graph into a set number of clusters or to satisfy some criteria for each cluster defined with respect to static node and/or edge weights. This differs significantly from the BN context and associated cost dependency between clusters.
@@ 1.3 Layout of Paper
We develop a novel algorithm called DCMAP (Dependent Cluster MAPping) that optimises BN cluster mappings given an arbitrary, non-negative scoring function which we refer to as a cost function in the presence of cost dependency. Such dependency arises in the BN context of using clustering to improve computational efficiency when inferring marginal distributions given evidence (Lauritzen and Spiegelhalter, 1988) and in model learning (Friedman et al., 1999; Lu et al., 2021). Noting as background the context of BN inference (Section 2.1) and introducing an illustrative example (Section 2.2), we first present an intuitive overview of DCMAP and its notation (Section 2.3), followed by its derivation from dynamic programming principles (Section 2.3.1 and 2.3.3). The algorithm is presented in detail in Section 3 and shown analytically to find all the optimal solutions such that an equally informed algorithm cannot do so in fewer iterations. An empirical demonstration is provided on the case study usage scenario of using clusters to minimise computation cost of inference for our complex systems seagrass DBN in Section 4. Reflections and suggestions for future work, including extension to other cost functions for optimisation such as entropy are provided in the Discussion (Section 5).
@@ 2.2 Illustrative Example
Intuitively, if the clustering cost function to be optimised (2) is dependent in some way on BN inference (3), then it will be affected by cost dependency due to the propagation of probabilities via parent-child arcs between clusters. To help illustrate this and the proposed algorithm, consider an example discrete BN DAG (Fig. 1). Assuming binary states for all nodes, the Conditional Probability Table (CPT) for each node with no parents (nodes $A$, $B$, $C$) is a $2\times 1$ matrix, e.g. $P(A)=[p_{a},p_{\bar{a}}]^{T}$ for states $A=a$ and $A=\bar{a}$. For $G$ and $F$ the CPT is a $2\times 2$ matrix, e.g. $P(F|A)=\begin{bmatrix}p_{af}&p_{\bar{a}f}\\ p_{a\bar{f}}&p_{\bar{a}\bar{f}}\end{bmatrix}$ where $p_{af}$ is the probability of state $F=f$ given state $A=a$. $P(D|A,B)$ is a $2\times 4$ matrix. Multiplication of CPTs (1) containing different nodes, i.e. variables, grows the dimensionality of the resultant matrix with associated exponential growth in computation; in the discrete case, the resultant matrix contains all possible combinations of node states. We ignore evidence terms for notational simplicity, noting that evidence can be easily incorporated using $\delta(X_{i})$ with node $X_{i}$ (1). Using (3), the posterior probability $P(F)$ is:
@@
$=$ $\sum\limits_{A}P(F|A)P(A)\sum\limits_{D,E,G}P(G|D,E)\sum\limits_{B}P(D|A,B)P(B)$
$\sum\limits_{C}P(E|C)P(C)$ (5)
Consider, as an example, a cost function based entirely on computational cost, and one approach where all nodes are in one cluster (4) where the product of conditional probabilities produces a matrix of $7$ dimensions, one for each node $A$ through $F$, with exponential growth in memory ($2^{7}$ joint probability values) and computation associated with multiplication ($608$ products) and marginalisation ($126$ sums). For example, $P(A)P(B)$ requires $4$ multiplications, and $P(A)P(B)P(D|A,B)$ requires $4+8=12$. $\sum_{A}P(F|A)$ requires $2$ summations.
@@ 2.3.1 Layers
We assume that $H$ is fully connected, since cluster mapping could be applied independently to each connected component. We use a simple breadth-first search (LaValle, 2006) to compute $l(\boldsymbol{X})$ in $O(n)$ time (Algorithm 1).
@@
This can be easily shown through contradiction since if there was an arc from $X_{1}$ to $X_{2}$, then $X_{1}$ is a parent of $X_{2}$ and Algorithm 1 would have updated $l(X_{1})$ to $l+1$ at the next iteration on line 7, which means they cannot be from the same layer. Hence, all nodes in the same layer are disjoint (see Fig. 1 for an example). Additionally, since there are no cycles in a DAG, Algorithm 1 is guaranteed to terminate after at most $n=|\boldsymbol{X}|$ iterations.

### v5

@@ 1 Introduction
Bayesian Networks (BNs) and related methods including Dynamics BNs (DBNs) are a form of graphical model that are represented as Directed Acyclic Graphs (DAGs), where relationships between nodes (factors) are shown as arcs and quantified with conditional probabilities [22]. They have been applied widely across domains including medicine, environment, economics and social sciences to provide explanatory decision support, integrating data and even expert knowledge [27, 11]. However, one of the key challenges associated with inference and learning of BNs and DBNs is computation, which is NP-hard [21, 19]. To this end, different approaches to clustering of BN nodes have been used in the context of supporting more computationally efficient inference of marginal distributions given evidence, and model learning, to update the structure of the graph and/or the conditional probabilities associated with each node [5, 15, 12, 18].
Optimal clustering of a BN seeks to minimise some cost or objective function associated with the mapping of BN nodes to clusters. Typically, each node is uniquely mapped to a cluster and the total cost is the sum of costs associated with each cluster. However, a key challenge is cost dependency. Consider a BN $H(\boldsymbol{X},\boldsymbol{A})$ with nodes $X_{i}\in\boldsymbol{X}$ and arcs (or edges) $\boldsymbol{A}$ that encode conditional dependence between parent and child nodes. Fundamentally, BN inference is underpinned by Bayes theorem where the joint probability distribution over all nodes $\boldsymbol{X}$ and given observations (i.e. evidence $\boldsymbol{E}$) is [15]:
@@
where $P(X_{j}|Par(X_{j}))$ is the conditional probability of node $X_{j}$ given its parents, and $\delta(X_{j})$ the evidence. For a discrete node with virtual evidence, $\delta(X_{j})=\{P(e|x_{j1}),\ldots,P(e|x_{jm})\}$ for binary event $e$ (observe this evidence or not) and $m$ node states $x_{j1},\ldots,x_{jm}$ [22]. The joint probability distribution arises from the multiplication of conditional probability distributions, each of which is potentially multi-dimensional, comprising a node (or variable) $X_{j}$ and parent nodes $Par(X_{j})$. Hence, intuitively: (i) clustering determines which conditional probability distributions are computed together, impacting computational complexity, and (ii) the computed outcomes of a given cluster are dependent on clusters connected by at least one arc via multiplication as per equation (1). This propagation of partial inference outcomes between clusters has been employed in approaches to BN and DBN inference such as Lauritzen and Spiegelhalter [15] and Wu et al. [26]. The dependency of the local cluster cost on the mapping of connected clusters is referred to as cost dependency, noting that the objective is to minimise the global cost which is a combination (typically additive) of local costs. However, finding optimal clusters (i.e. partitions) is NP-hard in general [4] and cost dependency exacerbates this complexity.
The clustering problem can be formalised as one of assigning a unique integer cluster label $u(X_{j})=k,k=1,\ldots,m$ to each node $X_{j}\in\boldsymbol{X}$ of BN $H(\boldsymbol{X},\boldsymbol{A})$; the total number of clusters $m$ is determined through optimisation. Let $G\bigl{(}H(\boldsymbol{X},\boldsymbol{A}),u(\boldsymbol{X})\bigr{)}$ be the total cost to be minimised through cluster mapping $u(\boldsymbol{X})$ and $G^{\ast}\bigl{(}H(\boldsymbol{X},\boldsymbol{A})\bigr{)}$ the optimal cost, and assume that the total cost is the sum of individual cluster costs $G_{k}$:
@@ 1.1 Case Study
We use as a case study a DBN of a seagrass ecosystem [26]. Seagrasses are a critical primary habitat for fish and many endangered species including the dugong and green turtle, and contribute USD\$1.9 trillion per annum through nutrient uptake and cycling. However, better management of human stressors such as dredging requires the modelling of cumulative effects arising from interactions between biological, ecological and environmental factors. The non-homogeneous DBN itself is complex and computationally intensive, and inference is infeasible within a standard (MCMC) framework for real-world decision support contexts. Furthermore, understanding areas of uncertainty in the system and designing data collection strategies to mitigate uncertainty is of utmost interest in managing such complex systems. This uncertainty can be characterised by entropy with clusters corresponding to scenarios of evidence (i.e. observed data) (Section 5.2). We focus on computation as a cost function $G$ to demonstrate how the proposed algorithm can find clusters to minimise the computational complexity of inference, and discuss how our algorithm could also be used to minimise other cost functions such as entropy.
@@ 1.2 Existing Work
Clustering of BNs has been employed in a wide range of contexts, evolving from a focus on computation of marginal distributions given evidence such as the seminal work of Lauritzen and Spiegelhalter [15] to predominantly causal discovery and graph learning [11, 18] in recent times. Although the approach of Lauritzen and Spiegelhalter [15] and related approaches such as Kschischang et al. [14] (see Daly et al. [5] for a review) employ clustering, they are based on a single mapping of nodes to clusters (referred to as cliques) using heuristics. They do not evaluate different cluster mappings on an outcome score. These challenges are addressed in more recent literature through three main approaches: (i) constraint based, that evaluate conditional independence structures i.e. if an arc exists or not, then its direction, (ii) scoring, often with heuristics to evaluate goodness of fit of a proposed network, and (iii) hybrid, combining constraints with heuristics [10] .
@@
A number of approaches use clusters as a structure to restrict or constrain the local search space in a graph learning framework such as the seminal work of Friedman et al. [7]. They define clusters in a similar vein to clique-trees [15] in a “restrict and maximise” framework to optimise a mutual-information based score of the learned network. Kojima et al. [12] advance this approach, effectively relaxing the cluster tree requirement in Friedman et al. [7] to a cluster graph to help address the issue of large clusters, which make local optimisation computationally infeasible. They used ancestral constraints to help limit the search space of clusters and the algorithm was shown to outperform greedy algorithms including Max-Min Hill Climbing (MMHC) and Hill Climbing (HC) with Tabu search. However, computational feasibility becomes a problem again when there are too many edges between clusters [12].
@@
Lu et al. [18] also adapt the work of Friedman et al. [7] for undirected graphs, using clusters based on proposed parents and children of a node $X$ to constrain the search in an iterative framework based on A*, a heuristic-informed search algorithm. They show empirically that for small networks, exhaustive and in some cases optimal methods such as A* and SAT (Boolean SATisfiability) outperform greedy methods such as PC (Peter and Clark), which can suffer from convergence to local optimal, and their adaptation of A* performs as well as greedy methods in larger networks. However, their approach is exponential in complexity with neighbourhood size (i.e. number of parents).
@@
Zhang et al. [28] present an alternate approach in which clusters are determined through agglomerative clustering using Pearson correlation as the distance metric; however, little is provided in terms of algorithm performance and its convergence characteristics. Huang and Zhou [10] and Gu and Zhou [9] both develop comprehensive approaches that use similar agglomerative clustering based on mutual information and correlation as the distance metric, respectively. The former learns skeletons (an undirected version of the graph) for each cluster using their variant of the PC algorithm and log-likelihood ratio tests of conditional independence. Gu and Zhou [9] use a score-based algorithm to learn partial DAGs for each cluster, which are then fused in the final step proposing edges between clusters. Huang and Zhou [10] combine clusters with automated parameter tuning for turning skeletons into partial then fully directed acyclic graphs and hybrid greedy initialisation, obtaining empirical improvements in computational efficiency and more accurate DAGs. However, in both cases, the performance is dependent on the quality and size of the clusters.
@@
Clearly, clustering has demonstrated utility to improve computational feasibility in learning BNs of varying degrees of optimality given a scoring function (i.e. cost function). However, all of the reviewed methods above focused on scores that changed given changes to the BN vis a vis inclusion or exclusion of edges, rather than scores that changed given different mappings of nodes to clusters. This paper focuses on the latter task of explicit optimisation with respect to clusters, to develop a more generalisable algorithm that works with arbitrary non-negative objective functions. Nevertheless, the literature does highlight the utility of dynamic programming and clustering for optimisation [7, 12, 18].
@@
Note that there also exists substantial literature in graph partitioning such as repeated graph bisection, Kernighan and Lin algorithm, spectral bisection and multilevel partitioning [4, 1]. Zheng et al. [29], for instance, demonstrate how dynamic programming can be used to optimally and efficiently cluster nodes in the absence of cost dependency, applied to convolutional neural networks. However, the problem context for these approaches typically involves partitioning a graph into a set number of clusters or to satisfy some criteria for each cluster defined with respect to static node and/or edge weights. This differs significantly from the BN context and associated cost dependency between clusters.
@@ 1.3 Layout of Paper
We develop a novel algorithm called DCMAP (Dependent Cluster MAPping) that optimises BN cluster mappings given an arbitrary, non-negative scoring function which we refer to as a cost function in the presence of cost dependency. Such dependency arises in the BN context of using clustering to improve computational efficiency when inferring marginal distributions given evidence [15] and in model learning [7, 18]. Noting as background the context of BN inference (Section 2.1) and introducing an illustrative example (Section 2.2), we first present an intuitive overview of DCMAP and its notation (Section 2.3), followed by its derivation from dynamic programming principles (Section 2.3.1 and 2.3.3). The algorithm is presented in detail in Section 3 and shown analytically to find all the optimal solutions such that an equally informed algorithm cannot do so in fewer iterations. An empirical demonstration is provided on the case study usage scenario of using clusters to minimise computation cost of inference for our complex systems seagrass DBN in Section 4. Reflections and suggestions for future work, including extension to other cost functions for optimisation such as entropy are provided in the Discussion (Section 5).
@@ 2.2 Illustrative Example
Intuitively, if the clustering cost function to be optimised (2) is dependent in some way on BN inference (3), then it will be affected by cost dependency due to the propagation of probabilities via parent-child arcs between clusters. To help illustrate this and the proposed algorithm, consider an example discrete BN DAG (Fig. 1). Assuming binary states for all nodes, the Conditional Probability Table (CPT) for each node with no parents (nodes $A$, $B$, $C$) is a $2\times 1$ matrix, e.g. $P(A)=[p_{a},p_{\bar{a}}]^{T}$ for states $A=a$ and $A=\bar{a}$. The CPT for $F$ is a $2\times 2$ matrix, i.e. $P(F|A)=\begin{bmatrix}p_{af}&p_{\bar{a}f}\\ p_{a\bar{f}}&p_{\bar{a}\bar{f}}\end{bmatrix}$ where $p_{af}$ is the probability of state $F=f$ given state $A=a$. $P(D|A,B)$ and $P(G|D,E)$ are $2\times 4$ matrices. Multiplication of CPTs (1) containing different nodes, i.e. variables, grows the dimensionality of the resultant matrix with associated exponential growth in computation; in the discrete case, the resultant matrix contains all possible combinations of node states. We ignore evidence terms for notational simplicity, noting that evidence can be easily incorporated using $\delta(X_{i})$ with node $X_{i}$ (1). Using (3), the posterior probability $P(F)$ is:
@@
$=$ $\sum\limits_{A}P(F|A)P(A)\left(\sum\limits_{D,E,G}P(G|D,E)\left(\sum\limits_{B}P(D|A,B)P(B)\left(\sum\limits_{C}P(E|C)P(C)\right)\right)\right)$ (5)
Consider, as an example, a cost function based entirely on computational cost, and one approach where all nodes are in one cluster (4) where the product of conditional probabilities produces a matrix of $7$ dimensions, one for each node $A$ through $G$, with exponential growth in memory ($2^{7}$ joint probability values) and computation associated with multiplication ($608$ products) and marginalisation ($126$ sums). For example, $P(A)P(B)$ requires $4$ multiplications, and $P(A)P(B)P(D|A,B)$ requires $4+8=12$. $\sum_{A}P(F|A)$ requires $2$ summations.
@@ 2.3.1 Layers
We assume that $H$ is fully connected, since cluster mapping could be applied independently to each connected component. We use a simple breadth-first search [16] to compute $l(\boldsymbol{X})$ in $O(n)$ time (Algorithm 1).
@@
This can be easily shown through contradiction since if there were an arc from $X_{1}$ to $X_{2}$, then $X_{1}$ is a parent of $X_{2}$ and Algorithm 1 would have updated $l(X_{1})$ to $l+1$ at the next iteration on line 7, which means they cannot be from the same layer. Hence, all nodes in the same layer are disjoint (see Fig. 1 for an example). Additionally, since there are no cycles in a DAG, Algorithm 1 is guaranteed to terminate after at most $n=|\boldsymbol{X}|$ iterations.